OpenStudy (anonymous):

Find the ratio in simplest form. Use "/" format for answer. 5 to 15

OpenStudy (anonymous):

divide by 5 1/3

OpenStudy (anonymous):

hey your back thanks. will u answer questions on this thread

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve for the variable. (simplify and use the "/" for the fraction bar)

OpenStudy (anonymous):

if that is 2/x=6/2 you cross multiply so 2*2=6*x 4=6x divide by 6 4/6=x

OpenStudy (anonymous):

ok, which 2 triangles are si,i; similar